# Proximal Interacting Particle Langevin Algorithms

### v2

@@ Abstract
We introduce a class of algorithms, termed Proximal Interacting Particle Langevin Algorithms (PIPLA), for inference and learning in latent variable models whose joint probability density is non-differentiable. Leveraging proximal Markov chain Monte Carlo (MCMC) techniques and the recently introduced interacting particle Langevin algorithm (IPLA), we propose several variants within the novel proximal IPLA family, tailored to the problem of estimating parameters in a non-differentiable statistical model. We prove nonasymptotic bounds for the parameter estimates produced by the different algorithms in the strongly log-concave setting and provide comprehensive numerical experiments on various models to demonstrate the effectiveness of the proposed methods. In particular, we demonstrate the utility of our family of algorithms on a toy hierarchical example where our assumptions can be checked, as well as for sparse Bayesian logistic regression, training of sparse Bayesian neural networks, and sparse matrix completion. Our theory and experiments together show that PIPLA family can be the de facto choice for parameter estimation problems in non-differentiable latent variable models.
@@ 1 Introduction
Latent variable models (LVMs) are a crucial class of probabilistic models which are widely used in machine learning and computational statistics in multiple applications such as image, audio, and text modelling as well as in the analysis of biological data [Bishop, 2006; Murphy, 2012]. LVMs have demonstrated great success at capturing (often interpretable) latent structure in data and are widely used across different scientific disciplines such as psychology and social sciences [Bollen, 2002; Marsh and Hau, 2007], ecology [Ovaskainen et al., 2016], epidemiology [Chavance et al., 2010; Muthén, 1992] and climate sciences [Christensen and Sain, 2012].
An LVM can be described as compactly as a parametrised joint probability distribution $p_{\theta}(x,y)$ (to be specified precisely later), where $\theta$ is a set of static parameters, $x$ denotes latent (unobserved, hidden, or missing) variables, and finally $y$ denotes (fixed) observed data. Given an LVM, there are two fundamental, intertwined statistical estimation tasks: (i) inference, which involves estimating the latent variables given the observed data and the model parameters through the computation of the posterior distribution $p_{\theta}(x|y)$, and (ii) learning, which involves estimating the model parameters $\theta$ given the observed data $y$ through the computation and maximisation of the marginal likelihood $p_{\theta}(y)$. These problems often need to be solved simultaneously. In particular, the learning problem is often termed maximum marginal likelihood estimation (MMLE) in the computational statistics literature, and the main challenge of learning in LVMs is that the marginal likelihood is often intractable to compute as we clarify below.
The marginal likelihood $p_{\theta}(y)$ (also called the model evidence [Bernardo and Smith, 2009]) in an LVM can be expressed as an integral, $p_{\theta}(y)=\int p_{\theta}(x,y)\mathrm{d}x$, over the latent variables. Hence the task of learning in an LVM can be compactly described as solving the following optimisation problem
@@
where $\Theta$ is the parameter space (which will be $\mathbb{R}^{d_{\theta}}$ in our setting throughout). A classical algorithm for this setting is the celebrated expectation-maximisation (EM) algorithm [Dempster et al., 1977], which was first proposed in the context of missing data. The EM algorithm is an iterative procedure consisting of two main steps. Given a parameter estimate $\theta_{k}$, the expectation step (henceforth E-step) computes the expected value of the log likelihood function $\log p_{\theta}(x,y)$ with respect to the current conditional distribution for the latent variables given the observed data $p_{\theta_{k}}(x|y)$, i.e., $Q(\theta,\theta_{k})=\mathbb{E}_{p_{\theta_{k}}(x|y)}[\log p_{\theta}(x,y)]$. The second step is a maximisation step (henceforth M-step) which consists of maximising the expectation computed in the E-step. The EM algorithm, when it can be implemented exactly, builds a sequence of parameter estimates $(\theta_{k})_{k\in\mathbb{N}}$ where $\theta_{k}\in\arg\max_{\theta}Q(\theta,\theta_{k-1})$, which monotonically increase the marginal likelihood, i.e., $\log p_{\theta_{k}}(y)\geq\log p_{\theta_{k-1}}(y)$ [Dempster et al., 1977].
The wide use of the EM algorithm is due to the fact that it can be implemented using approximations for both steps [Wei and Tanner, 1990; Meng and Rubin, 1993; Lange, 1995], leveraging significant advances in Monte Carlo methods and numerical optimisation techniques. In some classes of models, it is possible to analytically derive $p_{\theta}(x|y)$ for fixed $\theta$, which allows an exact (unbiased) Monte Carlo approximation for the E-step [Wei and Tanner, 1990]. The M-step is generally performed by numerical optimisation methods, most notably gradient descent [Meng and Rubin, 1993].
@@
When sampling from the posterior distribution is feasible, then the resulting algorithm is a form of stochastic approximation procedure with unbiased gradients, which is often amenable to analysis [Delyon et al., 1999]. However, in most interesting modern statistical models in machine learning, such posterior distributions are intractable. This issue can be circumvented by designing Markov chain Monte Carlo (MCMC) samplers for the E-step in practice. This has led to significant developments, where Markov kernels based on the unadjusted Langevin algorithm (ULA) [Roberts and Tweedie, 1996; Durmus and Moulines, 2017] have become a widespread choice in high dimensional setting thanks to their favourable theoretical properties [Durmus and Moulines, 2019; Vempala and Wibisono, 2019; De Bortoli et al., 2021; Chewi et al., 2022; Dalalyan, 2017]. Recently, Kuntz et al. [2023] explore an alternative approach based on Neal and Hinton [1998], where they exploit the fact that the EM algorithm is equivalent to performing coordinate descent of a free energy functional, whose minimum is the maximum likelihood estimate of the latent variable model. They propose several interacting particle algorithms to address the optimisation problem. This method has led to many subsequent works including Akyildiz et al. [2023]; Sharrock et al. [2024]; Lim et al. [2024]; Caprio et al. [2024]; Johnston et al. [2024]; Gruffaz et al. [2024] and ours.
@@
Contribution. In this work, we deal with models whose joint probability density is non-differentiable, exploiting the use of proximal methods [Combettes and Pesquet, 2011; Parikh and Boyd, 2014]. In the classical sampling case, this setting has been considered in a significant body of works, e.g., Pereyra [2016]; Durmus et al. [2018]; Atchadé et al. [2017]; Crucinio et al. [2023]; Diao et al. [2023]; Chen et al. [2022]; Bernton [2018]; Salim and Richtarik [2020]; Salim et al. [2019]; Lee et al. [2021] due to significant applications in machine learning, most notably in the use of non-differentiable regularisers. For example, this type of models naturally arises when including sparsity-inducing penalties, such as Laplace priors for regression problems or in the context of Bayesian neural networks [Williams, 1995; Yun et al., 2019], and total variation priors in image processing [Durmus et al., 2018].
Specifically, we summarise our contributions below.
@@
  • We develop the first proximal interacting particle Langevin algorithm family. Similar algorithms so far are investigated in the usual differentiable setting [Kuntz et al., 2023; Akyildiz et al., 2023; Johnston et al., 2024]. We extend these methods to the non-differentiable setting via the use of proximal techniques. Specifically, we propose two main algorithms, termed Moreau-Yosida interacting particle Langevin algorithm (MYIPLA) and proximal interacting particle gradient Langevin algorithm (PIPGLA).
@@
  • We provide theoretical analysis of the developed methods. We also provide, for comparisons, proximal extensions of the methods proposed in Kuntz et al. [2023], termed Moreau-Yosida particle gradient descent (MYPGD) and provide a full theoretical analysis under this setting.
@@
  • We implement our methods for a variety of examples within toy and realistic settings and demonstrate that the proximal interacting particle Langevin algorithms (PIPLA) family is a viable option for the MMLE problem - similar to the unadjusted Langevin algorithm (ULA) for sampling.
@@
This paper is organised as follows. In Section 2, we present the technical background necessary to develop our methods. In Section 3, we introduce the proposed algorithms—we provide their theoretical analysis in Section 4. Finally, in Section 5, we show comprehensive numerical experiments and we conclude with Section 6. The code is available in the GitHub repository: https://github.com/paulaoak/proximal-ipla.
1.1 Notation
@@
$W_{2}(\mu,\nu)=\left(\inf_{\gamma\in\mathbf{T}(\mu,\nu)}\int_{{}^{d}\times^{d}}\|x-y\|^{2}\mathrm{d}\gamma(x,y)\right)^{1/2}$
where $\mathbf{T}(\mu,\nu)=\{\gamma\in\mathcal{P}(^{d}\times^{d}):\gamma(A\times^{d})=\mu(A),\ \gamma(^{d}\times A)=\nu(A)\ \forall A\in\mathcal{B}(^{d})\}$ denotes the set of all transport plans between $\mu$ and $\nu$. In the following, we metrise $\mathcal{P}_{2}(^{d})$ with $W_{2}$.
@@ 2 Background
In this section, we introduce background material on Langevin dynamics and proximal methods. We also provide a brief overview of the recent literature on particle-based algorithms for MMLE in latent variable models.
@@ 2.1 Langevin Dynamics
At the core of our approach is the use of Langevin diffusion processes [Roberts and Tweedie, 1996], which are highly popular for building advanced sampling algorithms. To describe briefly, the Langevin stochastic differential equation (SDE) is given by
@@
where $U:\mathbb{R}^{d}\to\mathbb{R}$ is a continuously differentiable function and $(\mathbf{B}_{t})_{t\geq 0}$ is a $d$-dimensional Brownian motion [Pavliotis, 2014]. Under mild assumptions, (2) has a strong solution and $\pi(x)\propto e^{-U(x)}$ is the unique invariant distribution of the semigroup associated with the SDE [Pavliotis, 2014]. In most cases it is not possible to solve (2) analytically; however, we can resort to a discrete-time Euler-Maruyama approximation with step size $\gamma$ which gives the following Markov chain
$X_{n+1}=X_{n}-\gamma\nabla U(X_{n})+\sqrt{2\gamma}\xi_{n+1},$ (3)
where $(\xi_{n})_{n\in\mathbb{N}}$ is a sequence of i.i.d. $d$-dimensional standard Gaussian random variables. In the sampling literature, the algorithm (3) is known as the unadjusted Langevin algorithm (ULA) [Durmus and Moulines, 2019]. When $U$ is $\mu$-strongly convex and $L$-smooth (i.e. $\nabla U(x)$ is $L$-Lipschitz), ULA has very favourable properties. In particular, it can be shown that ULA converges to its own (biased) limit $\pi^{\gamma}$ exponentially fast, and that the asymptotic bias is of order $\gamma^{1/2}$ [Durmus and Moulines, 2019]. This makes ULA a favourable alternative to the Metropolis-adjusted Langevin algorithm (MALA) - which is asymptotically exact - because ULA does not require the computation of acceptance ratios and is cheaper in high dimensions [Durmus and Moulines, 2019].
@@ 2.2 MMLE with Langevin Dynamics
While Langevin dynamics and ULA, as described above, are a powerful tool for sampling from a given target measure $\pi$ such as a posterior distribution arising from a statistical model, powerful samplers are not enough in real applications as statistical models often have unknown parameters. In this case, it is necessary to combine the sampling methods with optimisation techniques to learn the parameters.
Given a statistical model $p_{\theta}(x,y):\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}\to\mathbb{R}$ with latent variables $x$ and parameters $\theta$ (assuming the observed data $y$ is fixed), a typical task for finding the optimal parameter is to maximise the marginal likelihood, i.e., solving the following problem
$\bar{\theta}_{\star}\in\arg\max_{\theta\in\Theta}p_{\theta}(y)=\arg\max_{\theta\in\Theta}\int_{{}^{d_{x}}}p_{\theta}(x,y)\mathrm{d}x.$ (4)
As mentioned in the introduction, classical techniques such as the EM algorithm are available for this task, however, they typically require many intermediate approximations which obscure their theoretical properties.
A recent approach is to build an extended stochastic dynamical system which can be run in the space $\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}$, with the aim of jointly solving the problem of latent variable sampling and parameters optimisation. In this avenue, Kuntz et al. [2023] first proposed a method termed particle gradient descent (PGD) which build on the observation made in Neal and Hinton [1998], that the EM algorithm can be expressed as a minimisation problem of the free-energy objective (see Appendix B.1 for details). By constructing the gradient flow w.r.t. this functional, Kuntz et al. [2023] arrive at the following system of SDEs
$\mathrm{d}\bm{\theta}_{t}^{N}$ $=-\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\bm{\theta}_{t}^{N},\mathbf{X}_{t}^{i,N})\mathrm{d}t,$
where $(\mathbf{B}_{t}^{i,N})_{t\geq 0}$ is a family of independent Brownian motions. The algorithm can be efficiently implemented using an Euler-Maruyama scheme. While the original paper does not contain a nonasymptotic analysis of the resulting discretisation, recently Caprio et al. [2024] have provided a nonasymptotic analysis of the algorithm under strong-convexity and Lipschitz continuity of $U$.
Inspired by this approach, Akyildiz et al. [2023] propose an interacting Langevin SDE
@@
where $(\mathbf{B}_{t}^{i,N})_{t\geq 0}$ is a family of independent Brownian motions. The main difference with Kuntz et al. [2023] is the noise term in the dynamics of $\theta$. This makes the system of SDEs a particular instance of a Langevin SDE, for which well-known tools and techniques can be leveraged for analysis. It is also argued that the added noise may help the method to escape local minima in nonconvex settings [Akyildiz et al., 2023]. Discretising (5), we obtain the interacting particle Langevin algorithm (IPLA)
$\theta_{n+1}$ $=\theta_{n}-\frac{\gamma}{N}\sum_{j=1}^{N}\nabla_{\theta}U(\theta_{n},X_{n}^{j,N})+\sqrt{\frac{2\gamma}{N}}\xi_{n+1}^{0,N},$
$X_{n+1}^{i,N}$ $=X_{n}^{i,N}-\gamma\nabla_{x}U(\theta_{n},X_{n}^{i,N})+\sqrt{2\gamma}\xi_{n+1}^{i,N},$
where $(\xi_{n}^{i,N})_{n\in\mathbb{N}}$ are i.i.d. standard Gaussian random variables of the corresponding dimension for all $i=0,1,\dots,N$. Under Lipschitz continuity and strong-convexity of $U$, IPLA has been shown to enjoy favourable convergence properties [Akyildiz et al., 2023].
Since both methods are proposed for differentiable $U$, both of these methods become inapplicable when $U$ is non-differentiable. In this paper, we propose a new class of algorithms, termed Proximal Interacting Particle Langevin Algorithms (PIPLA), which can be applied when the joint density of the model is non-differentiable.
2.3 Proximal Markov chain Monte Carlo
The last component we need to build parameter estimation methods in a non-differentiable setting is the well-known set of algorithms termed Proximal MCMC methods [Pereyra, 2016; Durmus et al., 2018; Klatzer et al., 2023; Habring et al., 2023; Ehrhardt et al., 2023; Durmus et al., 2019]. These methods leverage a powerful set of techniques termed proximal methods [Combettes and Pesquet, 2011; Parikh and Boyd, 2014] to handle the non-differentiability present in the target distribution.
The main idea behind proximal methods is to use proximity mappings of convex functions, instead of gradient mappings, to construct fixed point schemes and compute function minima [Combettes and Pesquet, 2011; Parikh and Boyd, 2014]. Constructing these iterative rules is made possible by the proximity operator, which is defined as follows.
Definition 2.1 (Proximity mappings).

### v3

@@ Abstract
We introduce a class of algorithms, termed proximal interacting particle Langevin algorithms (PIPLA), for inference and learning in latent variable models whose joint probability density is non-differentiable. Leveraging proximal Markov chain Monte Carlo techniques and interacting particle Langevin algorithms, we propose three algorithms tailored to the problem of estimating parameters in a non-differentiable statistical model. We prove nonasymptotic bounds for the parameter estimates produced by the different algorithms in the strongly log-concave setting and provide comprehensive numerical experiments on various models to demonstrate the effectiveness of the proposed methods. In particular, we demonstrate the utility of our family of algorithms for sparse Bayesian logistic regression, training of sparse Bayesian neural networks or neural networks with non-differentiable activation functions, image deblurring, and sparse matrix completion. Our theory and experiments together show that PIPLA family can be the de facto choice for parameter estimation problems in non-differentiable latent variable models.
@@ 1 Introduction
Latent variable models (LVMs) are a class of probabilistic models which are widely used in machine learning and computational statistics for various applications, such as image, audio, and text modelling as well as in the analysis of biological data [Bishop, 2006, Murphy, 2012]. LVMs have demonstrated great success at capturing (often interpretable) latent structure in data, which is crucial in different scientific disciplines such as psychology and social sciences [Bollen, 2002, Marsh and Hau, 2007], ecology [Ovaskainen et al., 2016], epidemiology [Chavance et al., 2010, Muthén, 1992] and climate sciences [Christensen and Sain, 2012].
An LVM can be described as compactly as a parametrised joint probability distribution $p_{\theta}(x,y)\propto e^{-U(\theta,x)}$, where $\theta$ is a set of static parameters, $x$ denotes latent (unobserved, hidden, or missing) variables, and finally $y$ denotes (fixed) observed data. Given an LVM, there are two fundamental, intertwined statistical estimation tasks that must be solved simultaneously: (i) inference, which involves estimating the latent variables given the observed data and the model parameters through the computation of the posterior distribution $p_{\theta}(x|y)$, and (ii) learning, which involves estimating the model parameters $\theta$ given the observed data $y$ through the computation and maximisation of the marginal likelihood $p_{\theta}(y)$. The learning problem is often termed maximum marginal likelihood estimation (MMLE) and the main challenge is that $p_{\theta}(y)$ is often intractable.
The marginal likelihood $p_{\theta}(y)$ (also called the model evidence [Bernardo and Smith, 2009]) in an LVM can be expressed as an integral, $p_{\theta}(y)=\int p_{\theta}(x,y)\mathrm{d}x$, over the latent variables. Hence the task of learning in an LVM can be formulated as solving the following optimisation problem
@@
where $\Theta$ is the parameter space (which will be $\mathbb{R}^{d_{\theta}}$ in our setting throughout). A classical algorithm for this setting is the celebrated expectation-maximisation (EM) algorithm [Dempster et al., 1977], which was first proposed in the context of missing data. The EM algorithm is an iterative procedure consisting of two main steps. Given a parameter estimate $\theta_{k}$, the expectation step (E-step) computes the expected value of the log likelihood function $\log p_{\theta}(x,y)$ with respect to the current conditional distribution for the latent variables given the observed data $p_{\theta_{k}}(x|y)$, i.e., $Q(\theta,\theta_{k})=\mathbb{E}_{p_{\theta_{k}}(x|y)}[\log p_{\theta}(x,y)]$. The second step is a maximisation step (M-step) which consists of maximising the expectation computed in the E-step. The EM algorithm, when it can be implemented exactly, builds a sequence of parameter estimates $(\theta_{k})_{k\in\mathbb{N}}$ where $\theta_{k}\in\arg\max_{\theta}Q(\theta,\theta_{k-1})$, which monotonically increase the marginal likelihood, i.e., $\log p_{\theta_{k}}(y)\geq\log p_{\theta_{k-1}}(y)$ [Dempster et al., 1977].
The wide use of the EM algorithm is due to the fact that it can be implemented using approximations for both steps [Wei and Tanner, 1990, Meng and Rubin, 1993, Lange, 1995], leveraging significant advances in Monte Carlo methods for the E-step and numerical optimisation techniques for the M-step. In particular, in most modern statistical models in machine learning, the posterior distribution, $p_{\theta}(x|y)$ for fixed $\theta$, is intractable, requiring an approximation for the E-step. One way to address this is by designing Markov chain Monte Carlo (MCMC) samplers. This approach has led to significant developments, where Markov kernels based on the unadjusted Langevin algorithm (ULA) [Roberts and Tweedie, 1996b, Durmus and Moulines, 2017] have become a widespread choice in high dimensional settings thanks to their favourable theoretical properties [Durmus and Moulines, 2019, Vempala and Wibisono, 2019, De Bortoli et al., 2021, Chewi et al., 2022, Dalalyan, 2017].
@@
Recently, Kuntz et al. [2023] explore an alternative approach for MMLE based on Neal and Hinton [1998], where they exploit the fact that the EM algorithm is equivalent to performing coordinate descent of a free energy functional, whose minimum is the maximum likelihood estimate of the latent variable model. They propose several interacting particle algorithms to address the optimisation problem. This method has led to subsequent works [Akyildiz et al., 2025, Sharrock et al., 2023, Lim et al., 2024, Caprio et al., 2024, Johnston et al., 2024, Gruffaz et al., 2024] including ours.
Contribution.
@@
This work focuses on LVMs whose joint probability density is non-differentiable by leveraging proximal methods [Combettes and Pesquet, 2011, Parikh and Boyd, 2014]. In the classical sampling case, this setting has been considered in a significant body of works, see, for example, Pereyra [2016], Atchadé et al. [2017], Durmus et al. [2018], Bernton [2018], Crucinio et al. [2025], Diao et al. [2023], Chen et al. [2022], Salim and Richtarik [2020], Salim et al. [2019], Lee et al. [2021], due to significant applications in machine learning, most notably in the use of non-differentiable regularisers. For example, this type of model naturally arises when including sparsity-inducing penalties, such as Laplace priors for regression problems or Bayesian neural networks [Williams, 1995, Yun et al., 2019], and total variation priors in image processing [Durmus et al., 2018]. They are also relevant for non-differentiable activation functions in neural networks. Specifically, we summarise our contributions below.
@@
  • We develop the first proximal interacting particle Langevin algorithm (PIPLA) family. Similar algorithms so far are investigated in the usual differentiable setting [Kuntz et al., 2023, Akyildiz et al., 2025, Johnston et al., 2024]. We extend these methods to the non-differentiable setting via the use of proximal techniques. Specifically, we propose two main algorithms, termed Moreau-Yosida interacting particle Langevin algorithm (MYIPLA) and proximal interacting particle gradient Langevin algorithm (PIPGLA).
@@
  • We present a theoretical analysis of our methods and, for comparison, of the proximal extensions we develop for the method in Kuntz et al. [2023], which we termed Moreau-Yosida particle gradient descent (MYPGD).
@@
  • We apply our methods to a variety of examples and demonstrate that the PIPLA family is a viable option for the MMLE problem in non-differentiable settings.
@@
This paper is organised as follows. Section 2 introduces the technical background necessary to develop our methods. In Sections 3 and 4, we present the proposed algorithms along with their theoretical analysis. Section 5 presents comprehensive numerical experiments before concluding.
Notation.
@@
In the following we adopt the convention that the solution to the continuous time solution is given by bold letters, whilst other processes (including the time discretisation) are not.
@@ 2 Background
We present the background and setting for our analysis.
@@ 2.1 Langevin Dynamics
At the core of our approach is the use of Langevin diffusion processes [Roberts and Tweedie, 1996b], which are widely used for building advanced sampling algorithms. The Langevin stochastic differential equation (SDE) is given by
@@
where $U:\mathbb{R}^{d}\to\mathbb{R}$ is a continuously differentiable function. Under mild assumptions, (2) admits a strong solution, and its associated semigroup has a unique invariant distribution given by $\pi(x)\propto e^{-U(x)}$ [Pavliotis, 2014]. In most cases, solving (2) analytically is not possible, however, we can resort to a discrete-time Euler-Maruyama approximation with step size $\gamma$ which gives the following Markov chain
$X_{n+1}=X_{n}-\gamma\nabla U(X_{n})+\sqrt{2\gamma}\xi_{n+1}.$ (3)
This algorithm, known as the unadjusted Langevin algorithm (ULA) [Durmus and Moulines, 2019], exhibits favourable properties when $U$ is $\mu$-strongly convex and $L$-smooth (i.e. $\nabla U(x)$ is $L$-Lipschitz). In particular, it converges exponentially fast to its biased limit $\pi^{\gamma}$ and the asymptotic bias is of order $\gamma^{1/2}$ [Durmus and Moulines, 2019].
@@ 2.2 MMLE with Langevin Dynamics
A recent approach to solve the MMLE problem in (1) is to build an extended stochastic dynamical system which can be run in the space $\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}$, with the aim of jointly solving the problem of latent variable sampling and parameters optimisation. Kuntz et al. [2023] first proposed a method termed particle gradient descent (PGD) which builds on the observation made in Neal and Hinton [1998], that the EM algorithm can be expressed as a minimisation problem of the free-energy objective (see Appendix B.1 for details). By constructing a gradient flow with respect to this functional, Kuntz et al. [2023] propose the following system of SDEs
$\mathrm{d}\mathbf{X}_{t}^{i,N}$ $=-\nabla_{x}U(\bm{\theta}_{t}^{N},\mathbf{X}_{t}^{i,N})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}^{i,N},\quad i=1,\dots,N,$
where $U(\theta,x)=-\log p_{\theta}(x,y)$ given fixed observations $y$. By introducing a noise term in the dynamics of $\theta$ (4), which may facilitate escaping local minima in non-convex settings, Akyildiz et al. [2025] propose an interacting Langevin SDE:
@@
$\mathrm{d}\mathbf{X}_{t}^{i,N}$ $=-\nabla_{x}U(\bm{\theta}_{t}^{N},\mathbf{X}_{t}^{i,N})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}^{i,N},\quad i=1,\dots,N.$
An Euler-Maruyama discretisation of (5), provides the interacting particle Langevin algorithm (IPLA). Under strong-convexity and smoothness of $U$, IPLA and PGD exhibit favourable convergence properties [Akyildiz et al., 2025, Caprio et al., 2024].
2.3 Proximal Methods
Proximal methods [Combettes and Pesquet, 2011, Parikh and Boyd, 2014] use proximity mappings of convex functions, instead of gradient mappings, to construct fixed point schemes and compute function minima. We now introduce some important definitions. Consider $U:\mathbb{R}^{d}\to\mathbb{R}$.
Definition 1 (Proximity mappings).